A train traveling 50 mph left a station 30 minutes before a second train running at 55 mph. How soon did the second train overtake the first?
If 50x represents the distance the slower train travels, then the faster train travels

Answer by stanbon(75887)

A train traveling 50 mph left a station 30 minutes before a second train running at 55 mph. How soon did the second train overtake the first?
---------------------------
1st train DATA:
rate = 50 mph ; time = x hrs ; distance = rt = 50x miles
------------------------------
2nd train DATA:
rate = 55 mph ; time = (x-(1/2)) hrs ; distance = rt = 55(x-(1/2)) miles
----------------------------
Equation:
distance = distance
50x = 55(x-(1/2)
50x = 55x - (55/2)
-5x = -55/2
x = 11/2
x = 5 1/2 hrs (time at which the 2nd train overtakes the 1st train)
==================================================================
Cheers,
Stan H.